OpenStudy (anonymous):

A rectangular tank with square ends has a length 1.5m greater than its width .The total length of all its edges are 20.4m. Find the dimensions of the square end of the box and find its volume

thomaster (thomaster):

so the total of all edges are 20.4m There are 4 lenght which are 1.5m greater than the width. The square ends have the same lenght and width. I think 20.4 - (4*1.5) = 14.4m <- this is the total length if it was a cube. A cube has 12 edge lines so 14.4 / 12 = 1.2m The height would be 1.2m, same for the height. The length would be 1.2+1.5 = 2.7m Don't know if i'm right :P not the best at math.

OpenStudy (anonymous):

just having ago at it myself .thanks for your reply.